# A Reinforcement Learning Method to Factual and Counterfactual Explanations for Session-based Recommendation

### v1

@@ 4.2. Recommendation Enhancement Module
We aim to enhance recommendation accuracy by fusing both factual explanations (FEs) and counterfactual explanations (CFEs). FEs serve as high-quality positive samples, while CFEs act as negative samples. By using these signals, we can fundamentally improve the SR model through contrastive learning, inspired by (Xie et al., 2022).
@@
Contrastive learning involves modifying the recommendation model’s loss function to integrate insights from these explanations. Accordingly, as discussed in (Oord et al., 2018), in SR we apply a contrastive loss function to differentiate positive samples from negative samples and maximize their separation, which is defined as:
@@ 5.1. Experimental Settings
Datasets: We adopt three datasets: Beauty, Baby, and Movielens-1Ms, each representing different consumer product categories from e-commerce platforms. For the Amazon datasets (Beauty and Baby222cseweb.ucsd.edu/ jmcauley/datasets.html.), the last session is assigned to the test set, with the remaining sessions used for training (Wu et al., 2023). Additionally, 10% of the training set is set aside as a validation set for model tuning. For MovieLens333grouplens.org/datasets/movielens/., 75% of the data is used for training, 15% for validation, and the remaining 10% serves as the test set to ensure balanced data partitioning. We split user interactions into sessions by day and exclude items with more than 100 interactions and sessions shorter than two interactions. This preprocessing strategy allows us to validate FCESR’s capability to generate explanations effectively across various session lengths. Table 2 shows the data statistics.
@@
#average training session length 6.64 5.65 9.47
@@
Baseline Models: Our work aims to provide explanations for existing SR models while leveraging the generated explanations as positive and negative samples for data augmentation to enhance their accuracy. Therefore, our baselines consist of two parts: explanation models for validating FCESR’s interpretability, and SR models for evaluating the effectiveness of our explanation-based data augmentation approach. For SR models, we consider several state-of-the-art backbones: GRU4REC (Hidasi et al., 2015) uses GRU to capture sequential interactions. NARM (Li et al., 2017) uses an attention mechanism to better capture evolving session interests. BERT4REC (Sun et al., 2019) employs bidirectional self-attention with a Cloze task for masked item prediction. SR-GNN (Wu et al., 2019) utilizes a gated GCN with self-attention for session representation. GCSAN (Xu et al., 2019) combines GNNs with self-attention to address both short- and long-term session dependencies. This diverse set includes RNN-based (GRU4REC), GNN-based (SR-GNN, GCSAN), attention-based (BERT4REC), and hybrid (NARM) approaches. We assess accuracy improvements by comparing each backbone with its FCESR-enhanced version (e.g., Base vs. FCESR), ensuring consistent inputs for a fair comparison.
Regarding explanation models, as the first work to propose both factual and counterfactual explanations in SR, we consider two baseline approaches for explanation generation: (1) Random Selection, a simple baseline that randomly samples items from the session as explanations. (2) CRVAE (Xu et al., 2021), a sequential recommendation explanation model that generates true explanations. Since no official implementation is available, we implemented it based on the pseudo-code provided in the original paper. CRVAE utilizes a VAE structure to perturb session data and identifies causally influential items using a sigmoid function.
@@
Dataset Metric GRU4REC NARM SRGNN GCSAN BERT4REC Base FCESR Improv. Base FCESR Improv. Base FCESR Improv. Base FCESR Improv. Base FCESR Improv. Baby HR@5 1.50 2.01 33.33%** 2.15 2.94 36.67%** 1.36 1.65 21.05%** 1.36 1.77 29.82%** 2.01 2.36 17.86%** HR@10 2.72 3.22 18.42%** 3.37 3.80 12.77%** 2.44 3.58 47.06%** 2.58 2.77 7.41%** 3.44 3.65 6.25%** HR@20 4.80 5.37 11.94%** 4.73 5.80 22.73%** 4.15 5.44 31.03%** 4.37 4.47 2.19%* 4.58 5.16 12.49%** NDCG@5 0.92 1.11 21.56%** 1.61 1.96 22.02%** 0.80 0.99 23.93%** 0.84 1.10 30.98%** 1.23 1.39 13.20%** NDCG@10 1.32 1.51 14.27%** 1.99 2.25 12.66%** 1.14 1.62 41.50%** 1.24 1.41 14.22%** 1.69 1.71 0.89% NDCG@20 1.84 2.04 10.79%** 2.33 2.75 17.66%** 1.57 2.08 32.49%** 1.68 1.83 9.09%** 1.97 2.15 8.95%** Beauty HR@5 19.15 19.18 0.17% 20.01 21.04 5.11%** 9.01 10.14 12.59%** 10.20 10.24 0.33% 12.07 14.74 22.10%** HR@10 22.01 22.21 0.91% 22.62 23.50 3.93%** 11.47 12.81 11.67%** 14.41 14.48 0.46% 14.41 17.28 19.91%** HR@20 25.15 25.22 0.27% 24.48 25.73 5.09%** 14.01 15.61 11.43%** 18.21 18.51 1.65%* 17.41 20.95 20.31%** NDCG@5 15.00 15.09 0.62% 16.93 17.11 1.10%* 7.03 7.89 12.23%** 8.67 8.96 4.32%** 8.97 11.63 29.61%** NDCG@10 15.96 16.10 0.88% 17.77 17.91 0.76% 7.83 8.75 11.72%** 8.24 8.52 3.43%** 9.72 12.46 28.15%** NDCG@20 16.75 16.85 0.60% 18.24 18.47 1.25%* 8.47 9.47 11.77%** 9.20 9.57 3.95%** 10.47 13.39 27.85%** MovieLens HR@5 7.74 8.68 12.14%** 10.46 11.69 11.77%** 6.75 7.46 10.44% 4.73 4.91 3.87%** 6.39 7.10 11.00%** HR@10 11.65 13.06 12.10%** 15.88 16.79 5.74%** 11.28 12.46 10.49%** 7.35 7.51 2.14%* 9.60 10.18 5.97%** HR@20 16.63 18.13 9.02%** 23.04 23.50 2.01%* 17.64 19.77 12.09%** 11.41 11.50 0.82% 14.58 15.08 3.46%** NDCG@5 5.11 6.46 26.42%** 7.19 7.32 1.85%* 4.28 4.90 14.53%** 3.13 3.29 5.24%** 4.19 4.59 9.44%** NDCG@10 6.37 7.14 12.09%** 8.93 9.55 6.94%** 5.74 6.51 13.41%** 3.97 4.13 3.98%** 5.23 5.43 3.87%** NDCG@20 7.63 8.83 15.73%** 10.73 11.92 11.08%** 7.33 8.34 13.81%** 4.99 5.60 5.13%** 6.48 6.59 1.80%*
@@
Regarding explanations, a good explanation should meet both factual and counterfactual conditions, meaning that the items included in the explanation should significantly influence the recommended results. We introduce the concepts of Probability of Sufficiency (PS) and Probability of Necessity (PN), derived from causal inference theory (Pearl et al., 2016) and used in GNN explanation task (Tan et al., 2022). They quantitatively assess the quality of explanations. In SR, we use them to evaluate whether the items generated by our method contribute to the final recommendation outcomes. Let $\mathbb{I}_{\mathcal{S}}^{K}$ be the set of items in the recommendation list. We define the PS and PN as:
@@
Hyper-parameter Setup: For each method, we tune hyper-parameters based on the performance on the validation sessions on each dataset. The dimensionalities of different layers in the neural network are uniformly set to 400 across all datasets. The batch size is consistently established at 256. For the Amazon datasets, the learning rate is set to 0.0001, while for the MovieLens dataset, it is set to 0.001. All models are trained using the Adam optimizer, and the RL discount factor is set to 0.99. Besides, during the fine-tuning process, we mainly focus on adjusting the learning rate and the coefficient $\lambda$ of the contrastive loss function to achieve optimal performance. The learning rate is varied in the range of [0.00001, 0.01], while $\lambda$ is tuned within the range of [0.1, 15] (Sun et al., 2020, 2022). Each method is executed five times to ensure statistical reliability, and average performance are reported. A paired t-test is conducted to compare each baseline with the FCESR_baseline to validate the significance of performance differences. Results are annotated with an asterisk (*) for $p\leq 0.05$ and a double asterisk (**) for $p\leq 0.01$.
Table 4. Comparative results regarding top-$1$ explanation on the three datasets. The best is boldfaced.
Models Baby Beauty MovieLens
@@
NARM-Random 97.79% 3.36% 6.49% 79.32% 16.78% 27.69% 83.09% 0.80% 1.58%
NARM-CRVAE 78.83% 33.10% 46.63% 77.35% 29.85% 43.07% 74.05% 24.77% 37.12%
NARM-FCESR 99.85% 28.89% 44.81% 99.27% 73.48% 84.45% 99.58% 13.45% 23.70%
SRGNN-Random 54.34% 8.51% 14.71% 86.21% 20.68% 33.36% 59.70% 4.35% 8.12%
SRGNN-CRVAE 79.46% 17.58% 28.79% 84.32% 15.92% 26.78% 79.31% 59.01% 67.67%
SRGNN-FCESR 99.46% 44.31% 61.31% 94.91% 53.52% 68.44% 89.49% 28.63% 43.38%
GRU4REC-Random 68.11% 28.67% 40.35% 64.41% 32.95% 43.59% 60.92% 27.16% 37.57%
GRU4REC-CRVAE 86.83% 24.67% 38.42% 76.24% 56.62% 64.98% 76.11% 41.89% 54.04%
GRU4REC-FCESR 98.63% 84.09% 90.78% 91.33% 80.48% 85.56% 99.56% 92.46% 95.88%
GCSAN-Random 48.65% 2.52% 4.79% 91.22% 4.05% 7.76% 58.64% 4.23% 7.88%
GCSAN-CRVAE 80.47% 17.60% 28.89% 74.30% 29.17% 41.89% 72.15% 38.46% 50.18%
GCSAN-FCESR 86.86% 29.08% 43.57% 99.11% 23.22% 37.63% 97.43% 32.49% 48.73%
BERT4REC-Random 46.24% 0.43% 0.84% 93.33% 2.76% 5.35% 71.17% 5.63% 10.43%
BERT4REC-CRVAE 86.14% 19.01% 31.14% 82.78% 18.20% 29.84% 89.90% 36.49% 51.91%
BERT4REC-FCESRC 87.65% 53.13% 66.16% 98.76% 19.92% 33.15% 99.74% 79.67% 88.58%
5.2. Recommendation Performance (RQ1)
The comparative results on five non-explainable SR methods are shown in Table 3. The following observations are noted. (1) FCESR substantially improves recommendation accuracy across all datasets and most evaluation metrics, showing remarkable enhancements compared to baseline models. (2) BERT4Rec and GRU4Rec exhibit significant performance improvements across all scenarios, whereas GCSAN’s improvements are relatively modest. This can be partially explained by the data in Table 4, which shows that BERT4Rec and GRU4Rec excel in generating effective explanations across all datasets. BERT4Rec, for example, achieves up to 79.67% effectiveness on MovieLens. This ability to produce high-quality explanations results in valuable samples that significantly boost accuracy after fine-tuning. In contrast, GCSAN’s lower explanation effectiveness (37.63% on Beauty, 48.73% on MovieLens) likely contributes to its smaller accuracy gains. These findings underscore that high-quality explanations not only enhance interpretability but also provide crucial training samples, thereby improving overall recommendation accuracy. This highlights the vital connection between explanation quality and recommendation effectiveness.
@@ 5.3. Explanation Performance (RQ2)
In FCESR, both factual and counterfactual reasoning are used to generate explanations. The results are presented in Table 4. We compare our results with randomly generated explanations and CRVAE for each method on different datasets and observe that FCESR successfully generates explanations for most of the recommended items, with following key findings: (1) FCESR consistently outperforms both Random and CRVAE baselines across most scenarios. (2) High PN values are observed across all models and datasets, likely due to the relative ease of generating counterfactual data by simply altering the predicted next item. (3) GCSAN exhibits lower PS values, particularly on BEAUTY (PS: 23.22%). This may be attributed to GCSAN’s strength in capturing local sequence features but its relative weakness in extracting higher-level global features, especially on complex, context-dependent datasets. (4) BERT4REC and GRU4REC show higher PS values, likely because of their superiority to capture long-term dependencies and contextual information in time series. BERT4REC’s bidirectional Transformers and GRU4REC’s RNN architecture excel at modeling complex sequential data, leading to more accurate and abundant explanations in challenging recommendation tasks.
@@
Baby FCESR (w/o) rank 99.58% 44.15% 61.18% 2.94
FCESR (w/o) spars 98.68% 43.42% 60.31% 5.61
FCESR (w/) fe 56.99% 68.24% 62.11% 5.63
FCESR 99.46% 44.31% 61.31% 2.93
MovieLens FCESR (w/o) rank 88.42% 52.13% 65.59% 3.77
FCESR (w/o) spars 74.53% 53.13% 62.04% 6.36
FCESR (w/) fe 63.72% 75.37% 69.06% 7.97
FCESR 89.49% 53.00% 66.57% 4.62

### v2

@@ 4.2. Recommendation Enhancement Module
We aim to enhance recommendation accuracy by fusing both factual explanations (FEs) and counterfactual explanations (CFEs). FEs serve as high-quality positive samples, while CFEs act as negative samples. By using these signals, we can fundamentally improve the SR model through contrastive learning, inspired by (Xie2022ContrastiveLF).
@@
Contrastive learning involves modifying the recommendation model’s loss function to integrate insights from these explanations. Accordingly, as discussed in (oord2018representation), in SR we apply a contrastive loss function to differentiate positive samples from negative samples and maximize their separation, which is defined as:
@@ 5.1. Experimental Settings
Datasets: We adopt three datasets: Beauty, Baby, and Movielens-1Ms, each representing different consumer product categories from e-commerce platforms. For the Amazon datasets (Beauty and Baby222cseweb.ucsd.edu/ jmcauley/datasets.html.), the last session is assigned to the test set, with the remaining sessions used for training (wu2023generic). Additionally, 10% of the training set is set aside as a validation set for model tuning. For MovieLens333grouplens.org/datasets/movielens/., 75% of the data is used for training, 15% for validation, and the remaining 10% serves as the test set to ensure balanced data partitioning. We split user interactions into sessions by day and exclude items with more than 100 interactions and sessions shorter than two interactions. This preprocessing strategy allows us to validate FCESR’s capability to generate explanations effectively across various session lengths. Table 2 shows the data statistics.
@@
#average training session length 6.64 5.66 9.47
@@
Dataset Metric GRU4REC NARM SRGNN GCSAN BERT4REC COCO-SBRS Base FCESR Improv. Base FCESR Improv. Base FCESR Improv. Base FCESR Improv. Base FCESR Improv. Base FCESR Improv. Baby HR@5 1.50 2.01 33.33%** 2.15 2.94 36.67%** 1.36 1.65 21.05%** 1.36 1.77 29.82%** 2.01 2.36 17.86%** 3.10 3.43 10.65%** HR@10 2.72 3.22 18.42%** 3.37 3.80 12.77%** 2.44 3.58 47.06%** 2.58 2.77 7.41%** 3.44 3.65 6.25%** 4.50 4.62 2.67% HR@20 4.80 5.37 11.94%** 4.73 5.80 22.73%** 4.15 5.44 31.03%** 4.37 4.47 2.19%* 4.58 5.16 12.49%** 6.20 6.25 0.81% NDCG@5 0.92 1.11 21.56%** 1.61 1.96 22.02%** 0.80 0.99 23.93%** 0.84 1.10 30.98%** 1.23 1.39 13.20%** 2.10 2.32 10.48%** NDCG@10 1.32 1.51 14.27%** 1.99 2.25 12.66%** 1.14 1.62 41.50%** 1.24 1.41 14.22%** 1.69 1.71 0.89% 2.60 2.75 5.77%** NDCG@20 1.84 2.04 10.79%** 2.33 2.75 17.66%** 1.57 2.08 32.49%** 1.68 1.83 9.09%** 1.97 2.15 8.95%** 3.00 3.13 4.33%** Beauty HR@5 19.15 19.18 0.17% 20.01 21.04 5.11%** 9.01 10.14 12.59%** 10.20 10.24 0.33% 12.07 14.74 22.10%** 9.49 9.79 3.21%** HR@10 22.01 22.21 0.91% 22.62 23.50 3.93%** 11.47 12.81 11.67%** 14.41 14.48 0.46% 14.41 17.28 19.91%** 16.13 16.43 1.83%** HR@20 25.15 25.22 0.27% 24.48 25.73 5.09%** 14.01 15.61 11.43%** 18.21 18.51 1.65%* 17.41 20.95 20.31%** 26.33 26.54 0.80%* NDCG@5 15.00 15.09 0.62% 16.93 17.11 1.10%* 7.03 7.89 12.23%** 8.67 8.96 4.32%** 8.97 11.63 29.61%** 5.64 5.78 2.51%** NDCG@10 15.96 16.10 0.88% 17.77 17.91 0.76% 7.83 8.75 11.72%** 8.24 8.52 3.43%** 9.72 12.46 28.15%** 7.78 7.81 0.45%* NDCG@20 16.75 16.85 0.60% 18.24 18.47 1.25%* 8.47 9.47 11.77%** 9.20 9.57 3.95%** 10.47 13.39 27.85%** 10.35 10.46 1.06%** MovieLens HR@5 7.74 8.68 12.14%** 10.46 11.69 11.77%** 6.75 7.46 10.44% 4.73 4.91 3.87%** 6.39 7.10 11.00%** 8.20 8.67 5.70%** HR@10 11.65 13.06 12.10%** 15.88 16.79 5.74%** 11.28 12.46 10.49%** 7.35 7.51 2.14%* 9.60 10.18 5.97%** 14.63 14.93 2.02%** HR@20 16.63 18.13 9.02%** 23.04 23.50 2.01%* 17.64 19.77 12.09%** 11.41 11.50 0.82% 14.58 15.08 3.46%** 24.46 24.78 1.33%** NDCG@5 5.11 6.46 26.42%** 7.19 7.32 1.85%* 4.28 4.90 14.53%** 3.13 3.29 5.24%** 4.19 4.59 9.44%** 5.00 5.21 4.22%** NDCG@10 6.37 7.14 12.09%** 8.93 9.55 6.94%** 5.74 6.51 13.41%** 3.97 4.13 3.98%** 5.23 5.43 3.87%** 7.06 7.23 2.43%** NDCG@20 7.63 8.83 15.73%** 10.73 11.92 11.08%** 7.33 8.34 13.81%** 4.99 5.60 5.13%** 6.48 6.59 1.80%* 9.52 9.87 3.64%**
Baseline Models: Our work aims to provide explanations for existing SR models while leveraging the generated explanations as positive and negative samples for data augmentation to enhance their accuracy. Therefore, our baselines consist of two parts: explanation models for validating FCESR’s interpretability, and SR models for evaluating the effectiveness of our explanation-based data augmentation approach. For SR models, we consider several state-of-the-art backbones: GRU4REC (hidasi2015session) uses GRU to capture sequential interactions. NARM (li2017neural) uses an attention mechanism to better capture evolving session interests. BERT4REC (sun2019bert4rec) employs bidirectional self-attention with a Cloze task for masked item prediction. SR-GNN (wu2019session) utilizes a gated GCN with self-attention for session representation. GCSAN (xu2019graph) combines GNNs with self-attention to address both short- and long-term session dependencies. COCO-SBRS (wang2023coco) models outer-session confounders (OSCs) to capture causal user intent.
Regarding explanation models, as the first work to propose both factual and counterfactual explanations in SR, we consider two baseline approaches for explanation generation: (1) Random Selection, a simple baseline that randomly samples items from the session as explanations. (2) CRVAE (xu2021learning), a sequential recommendation explanation model that generates true explanations utilizing a VAE structure. Since no official implementation is available, we implemented it based on the pseudo-code provided in the original paper. We also note that (cfe2024) does not explicitly generate counterfactual explanations but only outputs importance scores and rankings for items, and (li2024attention) does not provide reproducible source code. We attempted to reproduce their methods but were unable to obtain comparable implementations under the SR setting. Therefore, both methods are excluded from comparison.
@@
Regarding explanations, a good explanation should meet both factual and counterfactual conditions, meaning that the items included in the explanation should significantly influence the recommended results. We introduce the concepts of Probability of Sufficiency (PS) and Probability of Necessity (PN), derived from causal inference theory (pearl2016causal) and used in GNN explanation task (tan2022learning). They quantitatively assess the quality of explanations. In SR, we use them to evaluate whether the items generated by our method contribute to the final recommendation outcomes. Let $\mathbb{I}_{\mathcal{S}}^{K}$ be the set of items in the recommendation list. We define the PS and PN as:
@@
5.2. Recommendation Performance (RQ1)
The comparative results on five non-explainable SR methods are shown in Table 3. The following observations are noted. (1) FCESR substantially improves recommendation accuracy across all datasets and most evaluation metrics, showing remarkable enhancements compared to baseline models. (2) BERT4Rec and GRU4Rec exhibit significant performance improvements across all scenarios, whereas GCSAN’s improvements are relatively modest. This can be partially explained by the data in Table 4, which shows that BERT4Rec and GRU4Rec excel in generating effective explanations across all datasets. BERT4Rec, for example, achieves up to 79.67% effectiveness on MovieLens. This ability to produce high-quality explanations results in valuable samples that boost accuracy after fine-tuning. In contrast, GCSAN’s lower explanation effectiveness likely contributes to its smaller accuracy gains. These findings underscore that high-quality explanations not only enhance interpretability but also provide crucial training samples, thereby improving overall recommendation accuracy. This highlights the connection between explanation quality and recommendation effectiveness.
Table 4. Comparative results regarding top-11 explanation on the three datasets. The best is boldfaced.
Data Models Random CRVAE FCESR
@@ 5.2. Recommendation Performance (RQ1)
Baby NARM 99.95% 0.07% 0.14% 98.83% 23.10% 37.45% 99.91% 50.66% 67.23%
SRGNN 99.78% 0.21% 0.42% 99.46% 1.58% 3.11% 99.77% 2.12% 4.15%
GRU4REC 99.91% 0.04% 0.08% 98.83% 54.67% 70.40% 99.76% 96.51% 98.11%
GCSAN 99.59% 0.28% 0.56% 99.47% 0.60% 1.19% 99.86% 1.20% 2.37%
BERT4REC 99.92% 0.12% 0.24% 98.14% 19.01% 31.85% 98.26% 53.13% 68.97%
COCOSBRS 99.61% 9.40% 17.18% 98.90% 18.42% 31.06% 98.65% 32.88% 49.32%
Beauty NARM 99.94% 0.07% 0.14% 99.35% 29.85% 45.91% 99.27% 73.48% 84.45%
SRGNN 99.95% 0.03% 0.06% 98.32% 15.92% 27.40% 99.35% 18.39% 31.04%
GRU4REC 99.89% 0.04% 0.08% 99.24% 56.62% 72.10% 99.63% 98.51% 99.07%
GCSAN 99.95% 0.05% 0.10% 98.30% 6.17% 11.61% 99.11% 7.34% 13.67%
BERT4REC 99.90% 0.09% 0.18% 99.78% 18.20% 30.78% 98.76% 19.92% 33.15%
COCOSBRS 99.40% 0.07% 0.14% 99.93% 23.86% 38.52% 99.42% 31.27% 47.57%
MovieLens NARM 98.92% 0.11% 0.22% 98.05% 24.77% 39.55% 99.58% 33.45% 50.08%
SRGNN 99.77% 0.27% 0.54% 99.31% 19.01% 31.91% 99.49% 28.63% 44.46%
GRU4REC 99.15% 0.57% 1.13% 99.11% 41.89% 58.89% 99.56% 92.46% 95.88%
GCSAN 99.42% 0.09% 0.18% 98.15% 38.46% 55.26% 98.43% 32.49% 48.85%
BERT4REC 99.63% 0.07% 0.14% 98.90% 36.49% 53.30% 99.74% 79.67% 88.58%
COCOSBRS 98.37% 0.05% 0.10% 99.50% 33.10% 49.68% 99.46% 61.32% 75.87%
@@ 5.3. Explanation Performance (RQ2)
In FCESR, both factual and counterfactual reasoning are used to generate explanations. The results are presented in Table 4. We compare our results with randomly generated explanations and CRVAE for each method on different datasets and observe that FCESR successfully generates explanations for most of the recommended items, with following key findings: (1) FCESR consistently outperforms both Random and CRVAE baselines across most scenarios. (2) High PN values are observed across all models and datasets, likely due to the relative ease of generating counterfactual data by simply altering the predicted next item. (3) SRGNN and GCSAN exhibit lower PS values, particularly on the Baby dataset (PS: 2.12% and 1.2% respectively). This may be attributed to the reliance of SRGNN and GCSAN on consecutive transitions for graph construction, as removing items destroys this critical adjacency structure. (4) BERT4REC and GRU4REC show higher PS values, likely because of their superiority to capture long-term dependencies and contextual information in time series. BERT4REC’s bidirectional Transformers and GRU4REC’s RNN architecture excel at modeling complex sequential data, leading to more accurate and abundant explanations in challenging recommendation tasks.
@@
Beauty FCESR (w/o) spars 99.78% 98.79% 99.28% 6.62
FCESR (w/o) fe 98.25% 88.21% 92.96% 5.66
FCESR 99.63% 98.51% 99.07% 5.89
Baby FCESR (w/o) spars 98.68% 98.79% 98.73% 5.64
FCESR (w/o) fe 98.25% 88.23% 92.97% 5.63
FCESR 99.76% 96.51% 98.11% 5.58